OpenStudy (anonymous):

Solve. 5x^2 - 3x = (5x^2 - 7x) + (8x - 2) A. x = -2 B. x = 2 C. x = -1/2 D. x = 1/2

OpenStudy (anonymous):

5x^2 - 3x = 5x^2 + x - 2 -3x = x - 2 -4x = -2 x = 1/2